OpenStudy (anonymous):

For a quadratic equation in the form ax^2+bx+c=0, find the other root if one root is 8 and a=1 and b=-4

OpenStudy (anonymous):

x^2 -4x + c = 0 and (x - 8)(x + r) = 0 where r = other root sum of the roots = -b/a = --4/ 1 = 4 so 8 + r = 4 r = -4 other root = -4

OpenStudy (anonymous):

See, here a = 1, b = -2 and let c = c So you know: \[D = b^2 - 4ac\] D = 16 - 4c And roots can be found as: \[x = (-b \pm \sqrt{D})/2a\] It is given one of the root is 8: so, \[8 = (-(-4) + \sqrt{(16-4c)})/2\] \[8 = 2 + \sqrt{(4-c)}\] \[4 - c = 6^2\] c = -32 So, put this value in D.. \[D = 16 + 128 = 144\] So other root will be: \[x = (4 - \sqrt{D})/2\] Put the value of D and get the other value of x...

OpenStudy (anonymous):

So what if one root =1/2 a=-1 and c=6 ?

OpenStudy (anonymous):

Then use the product as Diajiones will explain it to you...

OpenStudy (anonymous):

@waterineyes the answer you gave me, is not right. so

OpenStudy (anonymous):

i dont know if you've done this in class yet by i used the following identities if A and B are the roots of the equation ax^2 + bx + c = 0 then A + B = -b/a and AB = c /a

OpenStudy (anonymous):

Huh?

OpenStudy (anonymous):

a = -1 c = -6 and 1 root is 1/2 then 1/2 * r = c/a = -6/-1 = 6 r/2 = 6 solve for r to get other root

OpenStudy (anonymous):

Did you solve it for x??? @Pamelaa23

OpenStudy (anonymous):

Im lost!

OpenStudy (anonymous):

can you solve this equation for ;